# Project Page: https://gyhdog99.github.io/projects/mocle/

### v4

@@ Abstract
Instruction tuning of Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, the diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflict for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve generalization capabilities of MoCLE for novel instructions. Extensive experiments on 11 zero-shot tasks demonstrate the effectiveness of MoCLE.
Keywords:
Instruction tuning Mixture of Experts Vision and language
@@ 1 Introduction
There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions [4, 48, 8], with instruction tuning emerging as a notably effective approach. This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks [41]. Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) [3, 66, 63, 10, 5], including LLaVA series [38, 37], MiniGPT-4 [73] and InstructBLIP [14]. They connect a trainable visual encoder with a frozen LLM and endow the LVLMs with the instruction-following ability by finetuning on diverse multi-modal instruction datasets.
It is observed in both the LLMs [51, 59, 13] and LVLMs [3, 69, 31] that the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks. However, the complexity of various instruction tasks brings difficulties for the model fine-tuning. Specifically, [60] find that under certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models, mainly attributed to the negative transfer phenomenon [68] during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading towards the sub-optimal experimental performance for these tasks.
@@
Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs, and thus, resulting in conflicts when sharing all the parameters as most studies adopt. In our preliminary study, we first split the instruction data into two disjoint subsets, including “cap” for image captioning and “vqa” for visual question answering. We then train InstructBLIP [14] using LoRA [21] on 3 kinds of data: “cap”, “vqa” and “full” (the full data) separately, to obtain 3 set of parameters (i.e., task experts). Following the held-out evaluation protocol [14], we evaluate these experts on the unseen datasets/tasks with the best expert. As shown in Figure 1, on 5 out of the 7 downstream tasks, the InstructBLIP instruction-tuned on all the data is outperformed by the task expert finetuned with only a subset of data. Among the 5 tasks, the Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”, revealing that instruction tuning on the similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks could hurt the generalization performance.
@@
The disjoint task expert above is a naïve solution to negative transfer, where we manually partition training tasks and train each expert individually. However, there are several limitations of this strategy: (1) The taxonomies such as “vqa” and “cap” require human expertise and become difficult to scale as the number of tasks grows. (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks, while some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1). In this regard, specialization and generalization of LVLMs becomes a dilemma.
@@
Our work aims to develop an automatic and practical partition strategy and a network architecture that can strike a balance between specialization and generalization. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning. Under the proposed MoCLE framework, we first cluster instructions of all the training data into several centroids via a pre-trained clustering model. In this way, similar tasks that could bring positive transfer to each other are automatically grouped into the same cluster, and different tasks that may cause conflict are separated (more justification for the usage of instruction clusters is detailed in Sec. 3.3). Then we construct several task experts with each concentrating on the specific cluster. Using the cluster as a condition, a router dispatches the input data to one of the specialized task experts and a universal expert that is shared among all the data. As we activate a specialized expert for a group of similar tasks, tasks that are less alike are learned via separate experts, mitigating task conflicts. Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
We validate effectiveness of MoCLE on the state-of-the-art InstructBLIP [14]. MoCLE achieves significant improvement on the held-out tasks [14] including the image captioning (e.g., Flickr30K [64]), the text-based visual question answering (e.g., TextVQA [56]), the knowledge-based visual question answering (e.g., ScienceQA [42]), and novel visual reasoning (e.g., Visual Spatial Reasoning [36]).
@@
The main contributions of this work contain three parts.
@@
    We identify the negative transfer phenomenon[40, 71] of task conflicts during the instruction tuning of LVLMs.
@@
    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate the task conflicts and enjoy the benefits of huge data training simultaneously.
@@
    Our proposed MoCLE achieves remarkable performance gains on the unseen held-out tasks compared with a strong baseline model InstructBLIP [14].
@@ 2 Related Work
2.0.1 Multi-task Instruction Tuning.
Instruction tuning, as elucidated in [51, 60], is a technique where a language model is finetuned over a collection of tasks and instruction templates that tells the model the intention of the task. The primary goal of this technique is to teach the model to understand the relationship between the instruction and the input/output pairs, so that the model could generalize to unseen task given novel instructions. Increasing the number of diverse instructions [51], the number of tasks [59, 13], and diversity of data [72] have been shown to improve the performance. However, [60] observe under certain model sizes, instruction tuning of language models fail to bring performance gains to unseen tasks when compared with untuned base models. They argue this is due to the fact all model capacity is used to learn the mixture of instruction tuning tasks. In this work, we aim to mitigate such issue in vision-language instruction tuning by the specialized experts.
2.0.2 Large Multi-modal Models.
To inherit strong reasoning and generalization abilities of LLMs and meanwhile understand visual inputs, recent works including Flamingo [2], BLIP-2 [28], LLaMA-adapter [67] connect LLMs to the visual modality with a frozen visual encoder and a trainable module, which is then finetuned with plain image-text pairs. To further improve instruction following abilities and response quality, LLaVA series [38, 37] finetune the whole LLM on the vision-language conversational data generated by GPT-4. MiniGPT-4 [73] performs training on ChatGPT-generated image captions instead of traditional caption data that are shorter and informational sparser. InstructBLIP [14] collects a wide range of academic VQA datasets and manually crafts 10 to 15 distinct instruction templates for each kind of task. In this paper, we primarily consider InstructBLIP as the base model.
2.0.3 Mixture of Experts.
MoE models [23, 24, 53] are considered as an effective way to increase the model capacity in terms of parameter size. Within MoEs, certain parts of the model are activated while the computation is kept the same or close to its dense counterpart. Recently, it has been thoroughly investigated in the field of computer vision[50, 71], natural language processing[27, 15] and multi-modal learning[47, 54, 12, 11, 34]. MoCE [71] tackles negative transfer in the self-supervised visual pre-training[6, 7] using a mixture of cluster experts. Different from MoCE requiring to be finetuned on downstream tasks, our MoCLE studies the zero-shot generalization to unseen tasks. Another line of works combines the MoE with adapters. AdaMix[58] randomly selects adapters during training and adopts the average of expert parameters during inference. Zadouri et al. [65] propose mixture of LoRAs for instruction tuning of LLMs.
Concurrently, MoE-LLaVA [34] sparsifies LVLMs via MoEs for more efficient model scaling. The most relevant to ours include the Octavius [12] and LLaVA-MoLE [11], which also address task conflicts in vision-language instruction tuning by applying the MoE architecture with LoRAs. However, they do not explore the issue of instruction generalization in the architecture designs. Instead, we design a novel Mixture of Cluster-conditional LoRA Experts together with the universal experts, which is the first work addressing both task conflicts and generalization with a unified architecture.
@@ 3 Methodology
In this section, we start with the formulation of LVLM instruction tuning and the analysis of the limitations of task experts. We then introduce the proposed MoCLE in detail, with the overall framework shown in Figure 2.
Figure 2: Overall pipeline of MoCLE. The left part is to cluster training data (a collection of vision-language tasks with diverse instructions) into several groups based on the instructions via $k$-means clustering algorithm. On the right is the MoCLE framework. $\mathbf{x}$ is the hidden representation of the sample (e.g., a sequence of tokens for images, instructions and the raw text inputs in the LLM) which is handled by 3 modules, including the universal expert, task expert, and the original linear module of the LLM. After that, the outputs of these modules are aggregated and passed to the next layer. Only experts, the cluster embedding and the gate parameters are updated during training, while parameters of the LLM are frozen. We insert such structure into the query and value projection in self-attention blocks of each layer by default.
@@ 3.1 Problem Formulation
Suppose that there is a set of datasets that are divided into held-in and held-out datasets [14], where a large vision-language model is first finetuned on the held-in and then evaluated on the held-out datasets in a zero-shot manner. To unify and diversify input-output formats and promotes in instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, a pair of text and image $(X_{\text{txt}},X_{\text{img}})$ from the corresponding dataset, where $T_{i}$ denotes the $i$-th template. For example, “Given the image, answer the question with no more than three words. {Question}” is a template for visual question answering tasks. The instruction is defined as $I\equiv T_{i}(X_{\text{txt}})$ that wraps text inputs using the template.
3.2 Limitation of Task Experts
Figure 1 indicates that allocating task experts based on the categories of training datasets brings positive transfer to similar downstream tasks. Nevertheless, this approach has several notable limitations. First, taxonomies such as “vqa” and “cap” require delicate human designs with expertise in the task performed. Another limitation is the challenge of selecting the appropriate expert for a newly introduced task when testing. Further, with tasks trained in an isolated manner, knowledge sharing among different tasks is prevented, compromising generalization to unseen tasks and instructions as some might benefit from multiple tasks. Hence, we propose MoCLE in the following sections to address these limitations.
@@ 3.3 Clustering Data by Instructions
The purposes of partitioning the training data are two-fold. We target at training a task expert within a collection of similar tasks so as to prevent task conflicts. Meanwhile, we expect novel tasks to be assigned to the proper experts based on their cluster without manual intervention. To achieve these, we should find a proper description for the task. Instructions serve as the foundation for LVLMs because they encapsulate the commands and intentions that guide responses of the model. The instructions $I$ in LVLMs consist of templates and the raw inputs. By understanding the templates in the instructions, the model can discern the type of task it needs to perform. The raw inputs further infuse this framework with the substantive content that the model must process and respond to. Therefore, by conducting clustering on the instructions, we can partition training data without human labor into several clusters, each containing a set of similar tasks, which is further verified through experiments in Sec. 4.5. Moreover, even when novel instruction tasks come, we can obtain a cluster index for them using the pre-trained clustering model without human intervention.
Formally, denoting $\mathcal{E}(\cdot)$ as a pre-trained sentence encoder, $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ is the sentence representation of an instruction $I_{i}$, we use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ denotes the set of instructions assigned to the $j$-th cluster and $\mathbf{c}_{j}$ denotes the centroid of the $j$-th cluster. In each iteration of the $k$-means clustering algorithm, each instruction is assigned to the nearest centroid with all centroids updated as the average of the instruction representations in the corresponding cluster.
@@ 3.4 Mixture of Cluster-conditional LoRA Experts
In addition to considerations at the data level, we also suggest an architectural design to tackle the issue of negative transfer. We propose Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data. Specifically, denoting $E$ as the number of experts, we introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$ that, given any input $\mathbf{x}_{i}$, determines the experts to which the input is routed. The gate vector is calculated as,
@@
where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets others to zero and $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all the layers, denotes the learnable embedding of the cluster that $\boldsymbol{x}$ belongs to. This is the key for the model to choose proper task experts for the input data. To endow the clustering embedding with task information, we initialize it to be the centroid of the corresponding cluster. $\mathbf{W}_{\text{gate}}$ denotes trainable weights of the linear gate, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ denotes a noise term that is designed to add some randomness to the expert choosing process, which encourages MoCLE to explore multiple combinations of experts during training. $\tau$ denotes a temperature hyperparamter. The output $\mathbf{y}_{i}$ is then computed as a sum of the weighted output of each expert and the original LLM linear layer [21] on the input $\mathbf{x}_{i}$ as,
@@
where $\mathbf{W}_{0}$ denotes the pre-trained linear layer of LVLMs, $\mathbf{W}_{e}$ is the linear projection weights of the $e$-th LoRA expert, and $G_{e}$, the $e$-th entry in $\mathbf{G}$, indicates the contribution of the $e$-th expert. In this formulation, the model learns to choose the expert that is suitable for the data by examining its cluster. We will show in Sec. 4.4.3 that the condition of clusters in expert routing is indispensable for task generalization with significant superiority over typical routing conditions including sentence embeddings and token representations.

### v5

# Mixture of Cluster-Conditional LoRA Experts for Vision-Language Instruction Tuning

@@ Abstract
Instruction tuning of Large Vision-language Models (LVLMs) has revolutionized the development of versatile models with zero-shot generalization across a wide range of downstream vision-language tasks. However, the diversity of training tasks of different sources and formats would lead to inevitable task conflicts, where different tasks conflict for the same set of model parameters, resulting in sub-optimal instruction-following abilities. To address that, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel Mixture of Experts (MoE) architecture designed to activate the task-customized model parameters based on the instruction clusters. A separate universal expert is further incorporated to improve generalization capabilities of MoCLE for novel instructions. Extensive experiments on InstructBLIP and LLaVA demonstrate the effectiveness of MoCLE.
Mixture of Cluster-Conditional LoRA Experts
for Vision-Language Instruction Tuning
Yunhao Gou1,2*, Zhili Liu2,3*, Kai Chen2*, Lanqing Hong3, Hang Xu3, Aoxue Li3, Xin Jiang3, Zhenguo Li 3, Dit-Yan Yeung2, James T. Kwok2, Yu Zhang1,4, 1Southern University of Science and Technology, 2The Hong Kong University of Science and Technology, 3Huawei Noah’s Ark Lab, 4Peng Cheng Laboratory, Equal contribution.  Correspondence: yu.zhang.ust@gmail.com
@@ 1 Introduction
There has been a continuously increasing trend to develop intelligent assistants that can follow human instructions Brown et al. (2020); OpenAI (2022); Chen et al. (2023c), with instruction tuning emerging as a notably effective approach. This method leverages large-scale well-formatted instruction data to empower Large Language Models (LLMs) to execute various human instructions, showcasing their ability to generalize across novel unseen tasks Longpre et al. (2023). Likewise, efforts have been made to introduce similar capabilities to Large Vision-language Models (LVLMs) Bai et al. (2023); Zhang et al. (2023); Ye et al. (2023); Chen et al. (2023e, a), including LLaVA series Liu et al. (2023b, a), MiniGPT-4 Zhu et al. (2023) and InstructBLIP Dai et al. (2023).
It is observed that for both the LLMs Sanh et al. (2021); Wang et al. (2022); Chung et al. (2022) and LVLMs Bai et al. (2023); Zhao et al. (2023); Li et al. (2023b), the ability to generalize to novel unseen instructions necessitates multi-task instruction tuning, i.e., training on a diverse collection of instruction-following tasks. However, the complexity of various instruction tasks brings difficulties for model fine-tuning. Specifically, Wei et al. (2021) find that for certain model sizes, multi-task instruction tuning even fails to bring performance gains for zero-shot tasks compared to the original models. This is mainly attributed to the negative transfer phenomenon Zhang and Yang (2017) during multi-task instruction tuning, where the model struggles to optimize the losses of multiple conflicted tasks, leading to sub-optimal performance.
@@
Similarly, tasks for vision-language instruction tuning (e.g., visual question answering and image captioning) focus on different perspectives of LVLMs. This results in conflicts as most studies adopt sharing of all parameters. In our preliminary study, we split the instruction data into two disjoint subsets (“cap” for image captioning, and “vqa” for visual question answering). We then train InstructBLIP Dai et al. (2023) using LoRA Hu et al. (2021) on three data sets (“cap”, “vqa” and the full data “full”) to obtain three sets of parameters (i.e., task experts). Following the held-out evaluation protocol Dai et al. (2023), we evaluate these experts on the unseen datasets/tasks with the best expert. As shown in Figure 1, on 5 out of the 7 downstream tasks, the InstructBLIP instruction-tuned on all the data is outperformed by the task expert finetuned with only a subset of data. Among the 5 tasks, Flickr30k belongs to “cap”, and SciQA, GQA and IconQA belong to “vqa”. This shows that instruction tuning on similar tasks brings positive transfer to downstream tasks, while training on the full data with dissimilar tasks can hurt generalization performance.
@@
The use of disjoint task experts above is a naïve solution to negative transfer, where we manually partition the training tasks and train each expert separately. However, it has several limitations: (1) The taxonomies such as “vqa” and “cap” require human expertise, and are difficult to scale as the number of tasks grows. (2) The ability to generalize to unseen tasks is inhibited, as we do not know which expert to choose for novel tasks, while some new tasks might benefit from multiple training tasks (e.g., VSR and TextVQA as in Figure 1). In this regard, specialization and generalization of LVLMs becomes a dilemma.
@@
This paper aims to develop an automatic and practical partition strategy and a network architecture that strikes a balance between specialization and generalization. In particular, we propose the Mixture of Cluster-conditional LoRA Experts (MoCLE) for vision-language instruction tuning. In this proposed framework, we first cluster instructions of all the training data into several clusters via a pre-trained clustering model. In this way, similar tasks that can bring positive transfer to each other are automatically grouped into the same cluster, while different tasks that may cause conflict are separated (more justifications for the use of instruction clusters are detailed in Sec. 3.2). Then we construct several task experts, with each focusing on a specific cluster. Using the cluster as condition, a router dispatches the input data to one of the specialized task experts and an universal expert that is shared among all data. As we activate a specialized expert for a group of similar tasks, tasks that are less similar are learned via separate experts, mitigating task conflicts. Meanwhile, since the universal expert trained on all tasks also contributes to the model outputs, we can enjoy generalization and specialization simultaneously.
We validate effectiveness of MoCLE on InstructBLIP Dai et al. (2023) and LLaVA-1.5 Liu et al. (2023a) and observe remarkable performance gains compared to dense models and other MoE baselinesChen et al. (2023f, 2024).
@@
The main contributions of this work contain the following three parts,
@@
    We identify the negative transfer phenomenon Liu et al. (2022b); Zhili et al. (2023) as tasks conflict during instruction tuning of LVLMs.
@@
    We propose Mixture of Cluster-conditional LoRA Experts (MoCLE), a novel parameter-efficient finetuning framework suitable for the vision-language instruction tuning, to mitigate task conflicts and enjoy the benefits of huge data training simultaneously.
@@
    Our proposed MoCLE achieves remarkable performance gains on held-in/out tasks compared to dense models and other MoE baselinesChen et al. (2023f, 2024).
@@ 2 Related Work
2.1 Multi-Task Instruction Tuning
Instruction tuning Sanh et al. (2021); Wei et al. (2021) fine-tunes a language model across tasks and instruction templates to convey task intentions. Its goal is to teach the model to understand relationships between instructions and input/output pairs, enabling generalization to unseen tasks with novel instructions. Increasing the number of instructions Sanh et al. (2021), tasks Wang et al. (2022); Chung et al. (2022), and data diversity Zhou et al. (2023) have shown to be effective in improving performance. However, Wei et al. (2021) find that for certain model sizes, instruction tuning fails to outperform untuned models on unseen tasks due to full capacity utilization for learning task mixtures. Our work addresses this in vision-language instruction tuning using specialized experts.
Figure 2: Overall pipeline of MoCLE.
2.2 Mixture of Experts (MoE)
MoE models Jacobs et al. (1991); Jordan and Jacobs (1993); Shazeer et al. (2017) are renowned for their ability to increase model capacity through parameter expansion. Recent research integrates MoE with adapters, exploring how pretrained adapters can be effectively combined Wu et al. (2024b), and how they enhance performance in both few-shot Huang et al. (2023) and zero-shot scenarios Jang et al. (2023); Muqeeth et al. (2024). Another line of research Chen et al. (2024); Wu et al. (2024a); Luo et al. (2024); Zadouri et al. (2023); Chen et al. (2023f) focuses on augmenting model capacity in a parameter-efficient manner. However, these approaches do not explicitly incorporate task/domain priors during the routing process, which might be limited when handling task conflicts. Two concurrent studies Liu et al. (2024); Li et al. (2024) incorporate domain information for expert routing. Unlike our approach, however, they do not address zero-shot generalization to unseen tasks.
@@ 3 Methodology
In this section, we start with the formulation of LVLM instruction tuning and an analysis of the limitations of task experts. We then introduce the proposed MoCLE. The overall framework is shown in Figure 2.
@@ 3.1 Problem Formulation
Suppose that there is a set of datasets that are divided into held-in and held-out datasets Dai et al. (2023). A large vision-language model is first finetuned on the held-in dataset, and then evaluated on the held-out dataset in a zero-shot manner. To unify and diversify input-output formats and promotes instruction tuning, several task templates $\{T_{i}\}$ are designed to wrap the raw inputs, which is a pair of text $X_{\text{txt}}$ and image $X_{\text{img}}$ from the dataset. For example, “Given the image, answer the question with no more than three words. {Question}” is a template for visual question answering tasks. The instruction is defined as $I\equiv T_{i}(X_{\text{txt}})$ that wraps text inputs using the template.
@@ 3.2 Clustering Data by Instructions
The purposes of partitioning the training data are two-fold. First, we hope to train a task expert with a collection of similar tasks so as to avoid task conflicts. Second, we expect novel tasks to be automatically assigned to the proper experts based on their cluster without manual intervention.
To achieve these goals, we conduct clustering on the instructions as they serve as the foundation for identifying different tasks. Formally, let $\mathcal{E}(\cdot)$ be a pre-trained sentence encoder, and $\mathbf{e}_{i}=\mathcal{E}(I_{i})$ be the sentence representation of an instruction $I_{i}$. We use the $k$-means clustering algorithm to group all instructions in the training datasets into $K$ clusters by iteratively minimizing $\sum_{j=1}^{K}\sum_{\mathbf{e}_{i}\in S_{j}}\|\mathbf{e}_{i}-\mathbf{c}_{j}\|^{2}$, where $S_{j}$ is the set of instructions assigned to the $j$th cluster, and $\mathbf{c}_{j}$ is the centroid of the $j$th cluster. In each $k$-means clustering iteration, each instruction is assigned to the nearest centroid with all centroids updated as the average of instruction representations in the corresponding cluster.
@@ 3.3 Mixture of Cluster-Conditional LoRA Experts
In addition to considerations at the data level, we also suggest an architectural design to tackle the issue of negative transfer. We propose the Mixture of Cluster-conditional LoRA Experts (MoCLE) that learns to activate the LoRA expert at each layer given the cluster of the data. Specifically, let $E$ as be the number of experts. We introduce a gate vector $\mathbf{G}\in\mathbb{R}^{E}$. Given an input $\mathbf{x}_{i}$, $\mathbf{G}$ determines the experts to which the input is routed. The gate vector is obtained as:
@@
where $\text{top}_{k}(\cdot)$ keeps the $k$ largest entries unchanged and sets the others to zero. $\mathbf{C}_{[\mathbf{x}]}$, which is shared among all layers, is the learnable embedding of the cluster that $\mathbf{x}$ belongs to. This is the key for the model to choose proper task experts for the input data. To endow the clustering embedding with task information, we initialize it to be the centroid of the corresponding cluster. Moreover, $\mathbf{W}_{\text{gate}}$ is the trainable weights of the linear gate, which is learned at each layer where the MoE block is inserted, $\boldsymbol{\epsilon}\sim N(0,\frac{1}{E})$ is a noise term that adds randomness to the expert choosing process (and encourages MoCLE to explore multiple combinations of experts during training [^1] We do not apply load balancing during training as we found it might distort task specialization.), and $\tau$ is a temperature hyperparameter. The output $\mathbf{y}_{i}$ is then computed as the sum of weighted outputs of the experts, and the original LLM linear layer Hu et al. (2021) on the input $\mathbf{x}_{i}$, as:
@@
where $\mathbf{W}_{0}$ is the pre-trained linear layer of LVLM, $\mathbf{W}_{e}$ is the linear projection weight of the $e$th LoRA expert, and $G_{e}$ (the $e$th entry in $\mathbf{G}$) indicates the contribution of the $e$th expert.
Models LLM Expert Params. # Experts # Clusters Rank Temperature Trainable Params. InstructBLIP Vicuna-7B q_proj, v_proj 4 + 1(universal) 64 8 0.05 Q-Former, LoRAs LLaVA-1.5 Vicuna-1.5-7B up_proj, down_proj 4 + 1(universal) 4 128 0.1 MLP connector, LoRAs
Table 1: Architecture details of MoCLE on different LVLMs. Note that experts are added to each layer of the transformer.